10g of hydrated calcium tetraoxosulphate(vi) [CaSO4.xH2O] weigh 7.90g after heating and cooling.calculate the emprical formular of the compound.


1
Expert's answer
2021-06-15T05:13:56-0400

delta m = 10 - 7.90 = 2.10 g

n = m / M

M (H2O) = 18 g/mol

M (CaSO4) = 136.14 g/mol

n (H2O) = 2.10 / 18 = 0.12 mol

n (CaSO4) = 7.90 / 136.14 = 0.06 mol

n (H2O) / n (CaSO4) = 0.12 / 0.06 = 2:1

The empirical formula of the compound CaSO4.xH2O is CaSO4.2H2O.